Accelerator Update for November 24, 2008 - December 1, 2008
Monday Nov 24
The day shift began with Operations monitoring the Tevatron (TeV) store 6589 (and the Recycler stash, with the Antiproton Source stacking, with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE experiment taking beam. NuMI was down due to a horn problem.
Tuesday Nov 25
The midnight shift began with Operations monitoring store 6590 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
The P1 transfer line beam permit tripped off and was reset. TeV experts changed the CDF collision point. Store 6590 was terminated and store 6591 (292.2E30) was established. Preaccelerator experts accessed the H- Source dome to adjust the extractor power supply. Linac experts replaced a modulator fan for Klystron RF station #3 (KRF3). Experts worked on the LCW mixing valve at MI-60; this inhibited stacking.
Wednesday Nov 26
The midnight shift began with Operations monitoring store 6591 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
MI RF station # 5 (MIRF5) tripped off and was bypassed. Store 6591 was terminated. TeV experts conducted studies. Operations established store 6593 (344.8E30). This store set a New Luminosity Record.
Thursday Nov 27
The midnight shift began with Operations monitoring store 6593 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
Operations terminated store 6593 and established store 6594 (293.7E30). The H- Source output dropped and began sparking. Operators switched to the I- Source. Booster vacuum got bad in period 6 and cascaded. Operators closed the vacuum valves. Linac RF station #2 tripped off. Linac and Booster experts worked on problems for most of the day shift.
Operations resumed sending beam through Linac and Booster at the beginning of the evening shift. An Accumulator power supply (A:QT) tripped off on a ground fault indication. KRF3 tripped off. Operations called in Pbar and Linac experts.
Friday Nov 28
The midnight shift began with Operations monitoring store 6594 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
During the midnight shift the MI-60 upper power supply tripped off. All the MI VCBs opened. Experts found and replaced a cable that had failed. Work on the MI-60 VCB continued.
On the evening shift Operations established store 6596 (267.6E30).
Saturday Nov 29
The midnight shift began with Operations monitoring store 6596 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
During the day shift Operations terminated store 6596 and established store 6597 (293.5E30). The I- Source output dripped to zero; a reboot fixed the problem. The Booster Gradient Magnet Power Supply (GMPS) tripped off. Operations contacted experts and operators reseated the supply’s control cards.
Sunday Nov 30
The midnight shift began with Operations monitoring store 6597 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
During the midnight shift the Booster Chopper stopped firing. Experts came in and replaced the “on” tube. Linac experts came in and replaced an extraction power supply. Operations terminated store 6597.
During the day shift Operations established store 6598 (328.6E30). D0 reported that they had lost a portion of their detector due to a power supply problem. The Booster GMPS tripped off due to a capacitor problem. Experts came in and removed the capacitor bank from the circuit. Pbar experts conducted studies.
During the evening shift Operations terminated store 6598. D0 accessed their collision hall and made repairs.
Monday Dec 1
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam.
During the midnight shift Operations established store 6601 (291.1E30). Cryo system experts at CHL reported that they had lost their cold box and its controls. Transfer line pressures were dropping. D0 and CDF put their detectors into standby.
The Plans for Monday
The plans for today are to stack and stash. The Cryo Coordinator reported that he needed to turn off the TeV cold compressors around the ring and it would not be back to normal operation until this afternoon. Operations will terminate the store before this happens.
Comments and Suggestions
|last modified 12/01/2008 email Fermilab|
|Security, Privacy, Legal|